OpenStudy (anonymous):

If cos=2/3 find csc; if 0°<0<90°

OpenStudy (jdoe0001):

\(\bf cos(\theta)=\cfrac{2}{3}\qquad 0^o<\theta<90^o \\ \quad \\ 0^o<\theta<90^o\Leftarrow\textit{another way of saying, "a" and "b" are positive} \\ \quad \\ cos(\theta)=\cfrac{2}{3}\implies \cfrac{adjacent}{hypotenuse}\implies \cfrac{a=2}{c=3}\qquad \textit{so what's "b"?} \\ \quad \\ c^2=a^2+b^2\implies \pm\sqrt{c^2-a^2}={\color{blue}{ b}}\qquad csc(\theta)=\cfrac{c}{{\color{blue}{ b}}}\) the pythagorean theorem doesn't tell us if the root value is +/-, so what's "b"? well we already know is in the I Quadrant, so is positive